OpenStudy (anonymous):

Help :( How many sides does the regular polygon have if each interior angle measure is four times the measure of each exterior angle measure?

Parth (parthkohli):

\[\textbf{HINT}\]The interior angle can be represented by \(\dfrac{180(n - 2)}{n}\) and the exterior angle can be represented by \(\dfrac{360}{n}\) . Now form an equation :)

OpenStudy (anonymous):

What numbers do I put in it :/? I was doing 4x+x= 180 :p

OpenStudy (anonymous):

I got 36 :p do i put 4 in it? :o

OpenStudy (anonymous):

I got 90 when I put 4 in.. and if its 90, ts a rectangle? @ParthKohli

OpenStudy (anonymous):

@Hero or someone check that please.. :p